# On Speeding Up Language Model Evaluation

### v1

@@ Abstract
Large language models (LLMs) currently dominate the field of natural language processing (NLP), representing the state-of-the-art across a diverse array of tasks. Developing a model of this nature, from training to inference, requires make numerous decisions which define a combinatorial search problem. For example, selecting the optimal pre-trained LLM, prompt, or hyperparameters to attain the best performance for a task often requires evaluating multiple candidates on an entire test set. This exhaustive evaluation can be time-consuming and costly, as both inference and metric computation with LLMs are resource-intensive. In this paper, we address the challenge of identifying the best method within a limited budget for evaluating methods on test examples. By leveraging the well-studied multi-armed bandit framework, which sequentially selects the next method-example pair to evaluate, our approach, combining multi-armed bandit algorithms with low-rank factorization, significantly reduces the required resources. Experiments show that our algorithms can identify the top-performing method using only 5-15% of the typically needed resources, resulting in an 85-95% reduction in cost.
@@ 1 Introduction
Figure 1: Comparison of cost for finding the best model or prompt between our proposed algorithms (UCB-E, UCB-E-LRF) and full evaluation on two datasets. Left: AlpacaEval [44] data as of May 20, 2024. The data includes 153 models evaluated on 805 instructions. The response on each instruction for each model is compared against a baseline response generated by GPT4-turbo, with GPT4-turbo being the LLM judge [66]. The monetary saving is calculated based on the inference cost of proprietary models among the 153 models and GPT4-turbo judging cost. Right: GSM8K [18] prompt engineering data. The data includes 205 zero-shot prompts for Mistral-7B generated by GPT4 evaluated on 784 questions (10%) of GSM8K training set. The time saving is calculated based on inference speed of Mistral-7B using the Huggingface library with one Nvidia A6000 GPU.
@@
Querying an LLM is resource-intensive, and therefore such extensive evaluations require significant investments in time, compute and financial resources. In Figure 1 Left, we show the estimated cost to fully evaluate (denoted as full evaluation) the 153 models officially included in AlpacaEval [44] as of May 20, 2024 to be almost 800 USD. In Figure 1 Right, we show that 78.2 Nvidia A6000 GPU hours are needed for evaluating 205 zero-shot prompts on 784 GSM8K [18] questions using Mistral-7B [32].
Despite the convention of exhaustive evaluations of all data points in a test set from a task, practitioners often only care about the overall performance rankings. Typically, the goal is to identify the top-performing methods or simply the best ones, discarding the lower-ranked alternatives. Therefore, while full-scale evaluation of each method on every data point is thorough, it may not be cost-effective when the goal is merely to identify the superior method.
@@
In this paper, we explore a limited-budget setting. The best method among a set of methods is the method that has the best average performance among a set of test examples. We want to identify this best method given a fixed budget for evaluating method-example pairs. A simple baseline is to evenly split the budget for each method. However, this can be very inefficient – as we observed in our experiments, for some datasets, we need to evaluate more than $90\%$ of all method-example pairs to predict the best method correctly with high probability. Intuitively, it is easier to recognize that low-performance methods are unlikely to be the best. Therefore, a more effective budget allocation strategy is to spend less on low-performing methods and more on promising ones. Multi-armed bandit algorithms enable this dynamic allocation by actively selecting the next method-example pair to evaluate based on the results of previous evaluations.
@@
We propose two active selection algorithms UCB-E and UCB-E-LRF. Our first algorithm is an extension of the classical UCB-E [2] to solve the multi-armed bandit problem. It estimates the upper confidence bound (UCB) to guide the selection of the method which is paired with a randomly chosen example for the next evaluation in order to efficiently estimate the best method. This algorithm enjoys a theoretical guarantee that the chance of selecting the best arm converges to 100% by an exponential decay of the number of evaluations. Our second algorithm, UCB-E-LRF, leverages the intrinsic low-rankness of the scoring matrices. We find that in practice, scoring matrices, where each row (column) is a method (an example) and the value is some metric score to measure how good the method is on the example, usually can be well-approximated by a low-rank matrix. Intuitively we can predict the remaining unobserved method-example pairs by the low-rank factorization (LRF) and prioritize evaluations of the pairs with large uncertainties in this prediction. By deploying this intuition in addition to UCB-E, UCE-E-LRF actively selects both the method and example to evaluate and it potentially improves the efficiency of the budget usage.
@@
We study the performance of these algorithms as well as common baselines in a number of practical settings. Through the empirical analysis, we find that our two active selection algorithms are much better than non-active baselines. In addition, we find for all algorithms, the harder settings, where the method set is large or the gaps between the top method and other methods are small, generally need more budget for evaluating the method-example pairs. More interestingly, we observe that the UCB-E works best for the easier settings; conversely, UCB-E-LRF has superiority over all other algorithms when the settings are harder.
@@ 2.1 Evaluation Workflow in Large Language Model Applications
A typical evaluation workflow in large language model (LLM) applications involves three steps: inference, scoring, and performance aggregation, the first two of which LLMs can play important roles.
@@
    Inference. Given a dataset such as one shown in Table 1 and an LLM-based method, the output from this method is generated through an LLM. Each method can be a distinct LLM for benchmarking different LLM performance, the same LLM with different prompts for prompt engineering [60], or other different configurations such as temperature, decoding strategies, etc for hyperparameter tuning.
@@
    Scoring. The outputs from different methods are scored with a scoring function i.e. metric. The scoring function can either be rule-based (exact string match, BLEU [50], ROUGE [46]), LLM-based (BERTScore [65], LLM judge [66]), or human-based i.e. user study. Depending on the task and dataset format, researchers have employed different types of scoring functions. In Table 1, we classify a selected group of commonly used datasets according to the task type and scoring function applied to them. We group LLM-based and human-based scoring together since they are usually considered alternatives to each other and have been shown to have relatively high correlation [44].
@@
    Performance aggregation. The performance of each method is aggregated across the dataset, typically with a simple average over all examples in the dataset.
Notice that LLMs can play an important role in both inference and scoring. However, LLMs are becoming increasingly large and many state-of-the-art LLMs for LLM-based scoring functions are black-box APIs [42]. Both inference and scoring in the evaluation workflow start to become massively resource-intensive.
Despite the intensive resources needed, in many practical scenarios, we are only interested in identifying the best method among all methods. For example, for prompt engineering and hyperparameter tuning, knowing which method (prompt / configuration) is the best is usually sufficient for next round of iteration. Intuitively, evaluating all methods on all examples on a dataset is excessive for this purpose and therefore we are interested in studying the identification of the best method with as few evaluation pairs as possible.
2.2 Notations and Problem Formulation
With the motivation above, we now define our notation and provide a formal problem formulation. Suppose we have a set of methods =$\mathcal{F}=\{f_{1},\ldots,f_{n}\}$ and a set of examples $\mathcal{X}=\{x_{1},\ldots,x_{m}\}$. Let $e:\mathcal{F}\times\mathcal{X}\to[0,1]$ denote our scoring function and without loss of generality, we assume $e(f,x)>e(f^{\prime},x)$ indicates $f$ has better performance than $f^{\prime}$ on the example $x$. Suppose $E\in[0,1]^{n\times m}$ is the underlying scoring matrix for a given problem $(\mathcal{F},\mathcal{X},e)$ by defining $E_{i,j}:=e(f_{i},x_{j})$. The score of a given method $f_{i}$ is defined as $\mu_{i}=\frac{1}{m}\sum_{j=1}^{m}E_{i,j}$ and want to find the best method $i^{*}=\arg\max_{i}\mu_{i}.$
As motivated above, we are interested in the scenario where we have a limited budget of $T$ evaluations. The question is: how can we select these $T$ method-example pairs to maximize the chance of finding the best method $i^{*}$? Formally, we want to study the evaluation algorithm $\mathcal{A}$. The input is the evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$. In the evaluation algorithm $\mathcal{A}$, we can evaluate at most $T$ method-example pairs $(f_{i_{t}},x_{j_{t}})$ ($t=1,\cdots T$). The output is the predicted best method. Our goal is to maximize $\mathbb{P}_{\mathcal{A}}(\mathcal{A}(T,\mathcal{F},\mathcal{X})=i^{*})$, the probability of returning the best method $i^{*}$.
@@ 3 Algorithms
One simple baseline for designing the evaluation algorithm $\mathcal{A}$ is: for each method $f_{i}\in\mathcal{F}$, we uniformly sample $\lfloor T/n\rfloor$ examples $\mathcal{X}_{i,T}$ from $\mathcal{X}$, estimate the mean $\hat{\mu}_{i}=\frac{1}{\lfloor T/n\rfloor}\sum_{x\in\mathcal{X}_{i,T}}e(f_{i},x)$, and pick the method $f_{\hat{i^{*}}}$ with the highest estimated mean $\hat{\mu}_{\hat{i^{*}}}$ as the prediction for the best method. However, this algorithm can be very inefficient. We will see in the experiments that for some datasets, we need to evaluate at least $90\%$ of all method-example pairs to predict the best method correctly with high probability. Different from this simple baseline, two algorithms proposed in this section actively decide the method-example pair to evaluate next and each decision is made based on previous evaluations; Figure 2 illustrates this high level idea.
Figure 2: Active method-example pair selection: After LLM evaluated $t$ method-example pairs, we then call Algorithm $\mathcal{A}$ to select the next method-example pair. Then we query LLM for evaluating this pair and fill the scoring received from LLM into the scoring matrix. Algorithm $\mathcal{A}$ then updates its internal status prepared for the next method-example pair selection. This process is repeated $T$ times and, in the end, the algorithm $\mathcal{A}$ predicts the best method $f_{\hat{i}^{*}}$
@@ 3.1 Algorithm 1 – UCB-E
The algorithm.
Notice the simple baseline evenly distributes its total budget across the different methods $f_{i}$. This allocation of resources is the main limitation of the simple baseline. In order to distinguish the best method $f_{i^{*}}$ from other good methods $f_{i}$ (i.e. $\mu_{i^{*}}-\mu_{i}$ is small), we may need more than $\lfloor T/n\rfloor$ examples, while $\leq\lfloor T/n\rfloor$ examples are sufficient to distinguish $f_{i^{*}}$ from bad methods (i.e. $\mu_{i^{*}}-\mu_{i}$ is large).
We address the limitation above with our first algorithm UCB-E ($\mathcal{A}_{ucb-e}$; Algorithm 1), a simple extension of the classic multi-arm bandit algorithm UCB-E [2]. At every step $t$, we estimate the upper confidence bound of each method $f_{i}$ from the examples evaluated with that method and pick the method $f_{i_{t}}$ with the largest upper confidence bound. Then we uniformly sample one example $x_{j_{t}}$ that has not been evaluated with the method $f_{i_{t}}$, run the inference procedure for $(f_{i_{t}},x_{j_{t}})$ and send the result to the annotator for scoring.
@@
The UCB-E algorithm addresses the limitation of the simple baseline. The upper confidence bounds of bad methods (i.e. $\mu_{i^{*}}-\mu_{i}$ is large) will never exceed the upper confidence bound of $f_{i^{*}}$, as long as they are evaluated with a sufficient number of examples, and hence UCB-E will never select this method again and pay more attentions on more promising methods.
Algorithm 1 UCB-E ($\mathcal{A}_{ucb-e}(T,\mathcal{F},\mathcal{X};\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$ and a set of examples $\mathcal{X}$, the uncertainty scaling $\eta$.
@@
1:  $\forall f_{i}\in\mathcal{F}$, the upper confidence bound $B_{i}=+\infty$, the set of evaluated examples $S_{i}=\{\}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$.
@@
3:     Select: Draw uniform randomly $i_{t}\in\arg\max_{i}B_{i}$; Draw uniform randomly $j_{t}\in[m]\backslash S_{i}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i_{t}},x_{j_{t}})$, send the result to the annotator, and receive $e(f_{i_{t}},x_{j_{t}})$; $E^{\rm obs}_{i_{t},j_{t}}\leftarrow e(f_{i_{t}},x_{j_{t}})$.
5:     Update: $S_{i_{t}}\leftarrow S_{i_{t}}\cup\{j_{t}\}$; $B_{i_{t}}\leftarrow\frac{\sum_{j\in S_{i_{t}}}E^{\rm obs}_{i_{t},j}}{|S_{i_{t}}|}+\sqrt{\frac{\eta}{|S_{i_{t}}|}}$.
@@
We state the theoretical guarantee of UCB-E in our context, which is a corollary of the theorem for UCB-E in the original multi-arm bandit setting.
Corollary 1 (Theoretical guarantee of UCB-E)
Define $H_{1}=\sum_{i=1,i\neq i^{*}}^{n}\frac{1}{(\mu_{i}-\mu_{i}^{*})^{2}}$ and suppose $\eta=\frac{25}{36}\frac{T-n}{H_{1}}$, $\mathbb{P}_{\mathcal{A}_{ucb-e}}(\mathcal{A}_{ucb-e}(T,\mathcal{F},\mathcal{X};\eta)=i^{*})\geq 1-2Tn\exp\left(-\frac{T-n}{18H_{1}}\right)$.
Comment. Given a problem set-up $(\mathcal{F},\mathcal{X},e)$, $H_{1}$ in the corollary indicates the hardness of the problem. The larger the gaps between the highest score and the remaining scores are, the higher the probability UCB-E will pick the best method given a fixed budget $T$. In our experiments, we will observe that the datasets with smaller $H_{1}$ tend to have a higher chance of predicting $i^{*}$ correctly given the same evaluation budget $T$.

### v2

@@ Abstract
Developing prompt-based methods with Large Language Models (LLMs) requires making numerous decisions, which give rise to a combinatorial search problem. For example, selecting the right pre-trained LLM, prompt, and hyperparameters to attain the best performance for a task typically necessitates evaluating an expoential number of candidates on large validation sets. This exhaustive evaluation can be time-consuming and costly, as both inference and evaluation of LLM-based approaches are resource-intensive. Worse, a lot of computation is wasted: Many hyper-parameter settings are non-competitive, and many samples from the validation set are highly correlated - providing little or no new information. So, if the goal is to identify the best method, it can be done far more efficiently if the validation samples and methods are selected adaptively. In this paper, we propose a novel method to address this challenge. We lean on low-rank matrix factorization to fill in missing evaluations and on multi-armed bandits to sequentially identify the next (method, validation sample)-pair to evaluate. We carefully assess the efficacy of our approach on several competitive benchmark problems and show that it can identify the top-performing method using only 5-15% of the typically needed resources—resulting in a staggering 85-95% LLM cost savings.
@@ 1 Introduction
Figure 1: Comparison of cost for finding the best model or prompt between our proposed algorithms (UCB-E, UCB-E-LRF) and Full Evaluation on two datasets. Left: AlpacaEval [44] data as of May 20, 2024. The data includes 153 models evaluated on 805 instructions. The response on each instruction for each model is compared against a baseline response generated by GPT4-turbo, with GPT4-turbo being the LLM judge [66]. The monetary saving is calculated based on the inference cost of proprietary models among the 153 models and GPT4-turbo judging cost. Right: GSM8K [18] prompt engineering data. The data includes 205 zero-shot prompts for Mistral-7B generated by GPT4 evaluated on 784 questions (10%) of GSM8K training set. The time saving is calculated based on inference speed of Mistral-7B using the Huggingface library with one Nvidia A6000 GPU. The overhead computation time for UCB-E and UCB-E-LRF is 2.4 and 142.6 seconds, respectively.
@@
Querying an LLM is resource-intensive, and therefore such extensive evaluations require significant investments in time, compute and financial resources. Figure 1 (left) shows the estimated cost of a Full Evaluation of the 153 models officially included in AlpacaEval [44] as of May 20, 2024 to be almost 800 USD. Similarly, 78 Nvidia A6000 GPU hours are needed for evaluating 205 zero-shot prompts on 784 GSM8K [18] questions using Mistral-7B [32], Figure 1 (right).
Although it is common practice to exhaustively evaluate all validation samples of a given task for each method, practitioners often only care about the final aggregated performance rankings. In other words, the goal is to identify the top-performing methods, discarding the lower-ranked alternatives. Consequently, evaluating on all validation samples is often unnecessarily costly and wasteful.
@@
In this paper, we explore a limited-budget setting. Given a set of methods we aim to identify the one that has the best average performance among a set of validation examples — given a fixed evaluation budget. A simple baseline is to evenly split the budget for each method. However, this can be very inefficient as it would spend a lot of the budget on obvious non-performers and too little effort on teasing out what the best methods are. Intuitively, methods that consistently underperform other methods are unlikely to end up as the best. Therefore, a more efficient budget allocation strategy is to spend less on low-performing methods and more on promising ones. Multi-armed bandit algorithms enable this dynamic allocation by actively selecting the next method-example pair to evaluate based on the results of previous evaluations.
@@
We propose two active selection algorithms UCB-E and UCB-E-LRF. Our first algorithm is an extension of the classical UCB-E [2] to solve the multi-armed bandit problem. It estimates the upper confidence bound (UCB) to guide the selection of the method which is paired with a randomly chosen example for the next evaluation in order to efficiently estimate the best method. This algorithm enjoys a strong theoretical guarantee that the chance of selecting the best arm converges to 100% by an exponential decay of the number of evaluations. Our second algorithm, UCB-E-LRF, leverages the fact that in practice many validation samples are similar, and their results are highly correlated. In other words, if a method performs well on one sample, it will predictably also perform well on similar samples. We can therefore speculatively fill in those results and evaluate the method on a sample that is less correlated. More formally, imagine we have $n$ methods and $m$ data points and store the evaluation results in a $(n\times m)$ matrix $E$. We can leverage the fact that this matrix $E$ is often low-rank, or can be approximated well by a low-rank matrix; evident from Table 3 and Figure 6 in Appendix A. We therefore regard the evaluation process as a matrix completion problem, where we predict the remaining unobserved $E_{i,j}$ entries by low-rank factorization (LRF) and prioritize evaluations of entries with large prediction uncertainties. By deploying this approach in addition to UCB-E, UCE-E-LRF actively selects both the method and example to evaluate, which leads to even bigger budget savings in settings with correlated validation samples.
@@
We evaluate the efficacy of these two algorithms, in addition to common non-active baselines, in a number of practical settings. Our empirical analysis shows that our two active selection algorithms are much better than all baselines. As predicted, the benefits of active selection are larger if the top methods are separated by a bigger performance gap. Notably, we observe that UCB-E works best for such easier settings; and UCB-E-LRF shines in harder settings, when the performance differences between methods are more subtle.
@@ 2.1 Evaluation Workflow in Large Language Model Applications
A typical evaluation workflow in large language model (LLM) applications involves three steps: inference, scoring, and performance aggregation.
@@
    Inference. Given a dataset (see Table 1), the output is generated by a specific LLM-based method. Here, methods can differ based on their LLM architecture (architecture search), a prompt (for prompt engineering [60]), or specific decoding parameters (e.g. temperature, decoding strategies), or hyper-parameters.
@@
    Scoring. The outputs from different methods are scored with a scoring function i.e. metric. The scoring function can either be rule-based (exact string match, BLEU [50], ROUGE [46]), LLM-based (BERTScore [65], LLM judge [66]), or human-based. Depending on the task and dataset format, researchers have employed different types of scoring functions. In Table 1, we classify a selected group of commonly used datasets according to the task type and scoring function applied to them. We group LLM-based and human-based scoring together since they are usually considered alternatives to each other and have been shown to have relatively high correlation [44].
@@
    Performance aggregation. The performance of each method is aggregated across the dataset, often with a simple average over all examples in the dataset, although many alternative aggregations are possible.
Both, inference and scoring do often rely heavily on LLM-based scoring through black-box APIs [42]—inducing large costs and resource consumption. However, in many practical scenarios, we are only interested in identifying the best method among all methods. For example, for prompt engineering and hyperparameter tuning, knowing which method (prompt / configuration) is the best is usually sufficient for next round of iteration. Identifying exactly which method is ranked 84 out of 100 is irrelevant and frankly, a waste of resources. We therefore claim that evaluating all methods on all examples of a dataset is excessive for this purpose, which motivates the research question how to identify the best method with as few evaluations as necessary.
2.2 Notation and Problem Formulation
Suppose we have a set of examples $\mathcal{X}=\{x_{1},\ldots,x_{m}\}$ and a set of methods =$\mathcal{F}=\{f_{1},\ldots,f_{n}\}$, such that $f_{i}:\mathcal{X}\to\mathcal{Y}$. Let $e:\mathcal{Y}\to[0,1]$ denote our scoring function and without loss of generality, we assume $e(f(x))>e(f^{\prime}(x))$ indicates $f$ has better performance than $f^{\prime}$ on the example $x$. Suppose $E\in[0,1]^{n\times m}$ is the underlying scoring matrix for a given problem $(\mathcal{F},\mathcal{X},e)$ by defining $E_{ij}:=e(f_{i}(x_{j}))$. The score of a given method $f_{i}$ is defined as $\mu_{i}=\frac{1}{m}\sum_{j=1}^{m}E_{ij}$ and we want to find the best method $i^{*}=\arg\max_{i}\mu_{i}.$ Formally, given a finite evaluation budget $T$, we want to study the evaluation algorithm ${\mathcal{A}}(\mathcal{F},\mathcal{X},T)$ and maximize $\mathbb{P}_{\mathcal{A}}(\mathcal{A}(T,\mathcal{F},\mathcal{X})=i^{*})$, the probability of returning the best method $i^{*}$ within budget $T$.
We introduce the following additional notation to allow for a clear description of the algorithms. Let $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ denote the partially observed scoring matrix, and $O\in\{0,1\}^{n\times m}$ denote the observation matrix where $O_{ij}$ indicates whether the method-example pair $(f_{i},x_{j})$ has been observed. Finally let us denote our estimate of $E$ by $\hat{E}$.
@@ 3 Algorithms
One simple baseline for designing the evaluation algorithm $\mathcal{A}$ is: uniformly sample and evaluate $T$ examples from $\mathcal{X}$, for each method $f_{i}\!\in\!\mathcal{F}$, estimate the performance $\hat{\mu}_{i}=\frac{1}{\sum_{j=1}^{m}O_{ij}}\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}$, and pick the method $f_{\hat{i}^{*}}$ with the highest estimated mean $\hat{\mu}_{\hat{i}^{*}}$ as the prediction for the best method. This baseline ignores any information gathered in the process of sequentially evaluating examples, and, as expected, can be very inefficient. We will see in the experiments that for some datasets, this baseline needs to evaluate at least $90\%$ of all method-example pairs to predict the best method correctly with high probability. In the following section, we will introduce two adaptive algorithms that actively select the next method-example pair to evaluate based on previous results; Figure 2 illustrates this idea on a high level for prompt tuning. The methods in the figure are different prompts, and the examples are simple math problems.
Figure 2: Active method-example pair selection: After LLM evaluated $t$ method-example pairs, we then call Algorithm $\mathcal{A}$ to select the next method-example pair. Then we query LLM for evaluating this pair and fill the scoring received from LLM into the scoring matrix. Algorithm $\mathcal{A}$ then updates its internal status prepared for the next method-example pair selection. This process is repeated $T$ times and, in the end, the algorithm $\mathcal{A}$ predicts the best method $f_{\hat{i}^{*}}$.
@@ 3.1 Algorithm 1 – UCB-E
The main limitation of the simple baseline is that in expectation it samples $\lfloor T/n\rfloor$ examples per method; evenly distributing its total budget across all different methods $f_{i}$. In order to distinguish the best method $f_{i^{*}}$ from other strong methods $f_{i}$ (i.e. $\mu_{i^{*}}-\mu_{i}$ is small), we may need more than $\lfloor T/n\rfloor$ examples, while $\leq\lfloor T/n\rfloor$ examples are sufficient to distinguish $f_{i^{*}}$ from obviously bad methods (i.e. $\mu_{i^{*}}-\mu_{i}$ is large).
We address this limitation by adaptive selecting example $t$ based on our previous $t-1$ observations until out budget $T$ is expended. Our first algorithm UCB-E ($\mathcal{A}_{\rm ue}$; Algorithm 1) is a simple extension of the classic multi-arm bandit algorithm UCB-E [2]. For every method $f_{i}$, at every step $t$, we estimate the upper confidence bound,
$B_{i}=\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}+\sqrt{\frac{a}{\sum_{j=1}^{m}O_{i,j}}}.$
In the subsequent iteration we pick the method $f_{i}$ with the largest upper confidence bound (ties are resolved by uniform random sampling), and then sample an example $x_{j}$ uniformly at random from the set of examples not yet observed, that is the set $\{j\in[m]|O_{ij}=0\}$. We run the inference procedure for $f_{i}(x_{j})$ and score the result, obtaining $E^{obs}_{ij}=e(f_{i}(x_{j}))$.
@@
The UCB-E algorithm efficiently addresses the limitations of the simple baseline by minimizing evaluations on suboptimal methods. For a bad method (i.e., when the gap $\mu_{i^{*}}-\mu_{i}$ is large), only a few evaluations (typically proportional to $\frac{1}{(\mu_{i^{*}}-\mu_{i})^{2}}$) are needed to ensure that its upper confidence bound will never exceed $\mu_{i^{*}}$. Consequently, the UCB-E algorithm will not select this method again, thus saving evaluation budget and focusing on more promising methods.
Algorithm 1 UCB-E ($\mathcal{A}_{\rm ue}(T,\mathcal{F},\mathcal{X};a)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, exploration parameter $a$.
@@
1:  The upper confidence bounds $B:=\{+\infty\}^{n}$, the observation matrix $O:=\{0\}^{n\times m}$, the observed scoring matrix $E^{\rm obs}:=\{?\}^{n\times m}$.
@@
3:     Select: Draw uniformly at random $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{m}O_{kj})\neq m}B_{k}$; Draw uniformly at random $j\in\{k\in[m]|O_{ik}=0\}$.
4:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
5:     Update: $O_{i,j}\leftarrow 1$; $B_{i}\leftarrow\frac{\sum_{j=1}^{m}O_{ij}\cdot E^{\rm obs}_{ij}}{\sum_{j=1}^{m}O_{ij}}+\sqrt{\frac{a}{\sum_{j=1}^{m}O_{ij}}}$.
@@
We state the theoretical guarantee of UCB-E in our context, which is a corollary of the theorem for UCB-E in the original multi-arm bandit setting [2].
Corollary 1 (Lower bound on success probability of UCB-E)
Define $H_{1}=\sum_{i=1,i\neq i^{*}}^{n}\frac{1}{(\mu_{i}-\mu_{i}^{*})^{2}}$ and suppose $a=\frac{25}{36}\frac{T-n}{H_{1}}$, $\mathbb{P}_{\mathcal{A}_{\rm ue}}(\mathcal{A}_{\rm ue}(T,\mathcal{F},\mathcal{X};a)=i^{*})\geq 1-2Tn\exp\left(-\frac{T-n}{18H_{1}}\right)$.
Comment. $H_{1}$ in the corollary indicates the hardness of the problem specified by $(\mathcal{F},\mathcal{X},e)$. The larger the gaps between the highest and all other scores are, the higher the probability that UCB-E will pick the best method given a fixed budget $T$. In our experiments, we will observe that the datasets with smaller $H_{1}$ tend to have a higher chance of predicting $i^{*}$ correctly, given the same evaluation budget $T$.